Is 1 + x + [x /(1 )] /(1 ) an identity or equation? Explain by

2 − x = 1 − x

giving at least 2 example solutions.


1
Expert's answer
2021-02-26T05:40:30-0500

1 + x + [x /(1 )] /(1 ) This problem is not an equation nor identity as it has no sign "=". You can just simplify it. And it will be 1+2x.

2 − x = 1 − x This one is incorrect identity, as in the end is 2=1.

So, Solving an equation means finding the value or values for which the two expressions are equal.

2x-1=5 ; 2x=6; x=3.


... An identity is an equation which is always true, no matter what values are substituted. 2 x + 3 x = 5 x. is an identity because 2 x + 3 x will always equal regardless of the value of .